Session: Oct 2007 (session 3)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (orfei at Wed Oct 31 00:34:38 2007) Success :
Raining for almost all the experiment

Torun (Kaz Borkowski at Wed Oct 31 07:44:19 2007) Minor failures :
The first hour has not been recorded due to problems with Mark5A version 2007y222d00h (used during the previous experiment with the result that our FTP for the fringe test failed altogether). The version 2005y147d11h has been restored for the rest of this session.

Onsala (Michael Lindqvist at Wed Oct 31 07:54:46 2007) Success :
Very windy, close to the limit of 13 m/s. Does affect pointing at 6.7 GHz. Otherwise no known problems.

Westerbork (Tony Foley at Wed Oct 31 09:05:19 2007) Minor failures :
Single dish was used. Please change the coordinates in the vexfile accordingly. IFD readout sometimes WRONG (said ifd=alt,nor when it was ifd=nor,nor)

Noto (P. Cassaro at Wed Oct 31 10:54:34 2007) Success :
No known problems

Jodrell_Bank (Alastair Gunn at Fri Nov 2 11:54:00 2007) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Fri Nov 2 11:54:39 2007) Minor failures :
Noticed receiver carousel in wrong position and corrected at 15:04 UT. Data probably lost up to this point. No other known problems.

Effelsberg (A. Polatidis at Mon Nov 5 11:01:46 2007) Success :